OpenStudy (anonymous):

Which are partial products for 535 × 69? Choose all answers that are correct. A. 45 B. 180 C. 450 D. 1,800

OpenStudy (anonymous):

@swilson90

OpenStudy (anonymous):

@Goose.

OpenStudy (anonymous):

@Jacob902

OpenStudy (anonymous):

i swear its gotta be all of them..

OpenStudy (anonymous):

cant be

OpenStudy (anonymous):

guess i can try it

OpenStudy (anonymous):

blah , you could try but ...maybe look it up on another website?

OpenStudy (excalibur0126):

However many points the question is worth is how many answers there are.

OpenStudy (anonymous):

what do u mean?

OpenStudy (excalibur0126):

If this is online school, then it will say how many points the question is worth, and that's how many answers there are correct

OpenStudy (anonymous):

its worth 1 point..

OpenStudy (anonymous):

@CallMeKiki

OpenStudy (anonymous):

@a1234

OpenStudy (anonymous):

what the monkey is a "partial product" ??

OpenStudy (anonymous):

what? @satellite73

OpenStudy (anonymous):

A partial product is pretty much breaking down a multiplecation problem to the expanded form (Like we learned you can do with back in elementary school, 10*10=5*5*4)

OpenStudy (anonymous):

hmm i doubt it

OpenStudy (anonymous):

But there is a slight difference

OpenStudy (anonymous):

that is called "factoring"

OpenStudy (anonymous):

maybe like \(5\times 9=45\)?

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

It's a difference. You have to break the problem down according to the place each digit is in (ones place, tens place, hundreds place etc.) and they have to add up to get to the original number.

OpenStudy (anonymous):

539 x 69 would be (500 + 30 + 9) times 69

OpenStudy (anonymous):

you sure about that ^^

OpenStudy (anonymous):

that is 37,191

OpenStudy (anonymous):

(500 + 30 + 9) times (60 + 9)

OpenStudy (anonymous):

Now we have to find the partial products

OpenStudy (anonymous):

yes what are the partial products though.. O.O

OpenStudy (anonymous):

So this is what you do, With the numbers 60 and 9, you have to determine which combination of them will get you to the answer of 539*69

OpenStudy (anonymous):

You have to multiply either 500, 60, or 5 by 60 or 9 in a certain combination to get you to the same answer the original problem (539*69) would get you to.

OpenStudy (anonymous):

You can use any number more than once as long as it is not in the same problem.

OpenStudy (anonymous):

Let's try 9 first and see can we get 1,800 out of that to get a part of the answer (i.e a partial product)

OpenStudy (anonymous):

i guess it is too much trouble just to say "multiply" gotta make it more confusing

OpenStudy (anonymous):

5 and 9 makes 45 when you multiply them, so that could be a partial product, let's see if we can get the rest of the equation to verify

OpenStudy (anonymous):

500 times 60 is equal to 30k, so now we have 30,045

OpenStudy (anonymous):

500 times 9 is 4500, so now we have 34,545 out of 36,915

OpenStudy (anonymous):

30*60=1800, so now we have 36345 out of 36915

OpenStudy (anonymous):

So 1800 can be a partial product too, so now we have 45 and 1800 and partial products

OpenStudy (anonymous):

wait, sorry about the 45, that's impossible.

OpenStudy (anonymous):

every1 said 45 how is that impossible O.O

OpenStudy (anonymous):

500 + 30 + 9; 500 is not equal to 5.

OpenStudy (anonymous):

ohh

OpenStudy (anonymous):

450 looks like it's the final partial product.

OpenStudy (anonymous):

the answers are... 1800 and..180..??

OpenStudy (anonymous):

So 1800 and 450 seems to be the partial products

OpenStudy (anonymous):

Oh waiitttt 535, I thought that was a 9

OpenStudy (anonymous):

Okay, so 45, 1800 and 450

OpenStudy (anonymous):

On second thought, I can't verify 450, I don't even have any work for that one.

OpenStudy (anonymous):

x-x oh mah gosh